# Evidence of partonic collectivity in ultra-relativistic heavy-ion collisions with NCQ scaling of radial flow

### v1

@@ Abstract
We report the first observation of Number of Constituent Quark (NCQ) scaling of the radial flow observable $v_{0}(p_{\rm T})$ in relativistic heavy-ion collisions. Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV are investigated using the string-melting version of the AMPT model, while Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV are analyzed using published ALICE data, with calculations from PYTHIA8/Angantyr serving as a non-collective baseline. The $v_{0}(p_{\rm T})$ spectra from AMPT-SM exhibit clear collective signatures, including mass ordering at lower-$p_{\rm T}$ and a meson–baryon separation at intermediate-$p_{\rm T}$, which are absent in the baseline calculations. In central collisions, $v_{0}(p_{\rm T})/n_{q}$ follows robust NCQ scaling when expressed as a function of transverse kinetic energy per quark, $(m_{\rm T}-m_{0})/n_{q}$, while significant deviations emerge toward peripheral events, indicating a stronger manifestation of collectivity in central collisions. The scaling is more precise at RHIC than at LHC energies, consistent with earlier observations for elliptic flow $v_{2}$. These findings provide strong evidence that radial collectivity is established predominantly at the partonic stage of the fireball evolution.
keywords:
QGP , radial flow , NCQ scaling , transport model , PYTHIA8/Angantyr
journal: Physics Letters B
@@
[label1]organization=Department of Physics, Bodoland University, addressline=Rangalikhata, city=Kokrajhar, postcode=783370, state=Assam, country=India
[…]
[label2]organization=Department of Physics, Kokrajhar University, city=Kokrajhar, postcode=783370, state=Assam, country=India
1 Introduction
Ultra-relativistic nucleus-nucleus collisions provide access to hot and dense QCD matter, commonly known as the quark-gluon plasma (QGP), where quarks are no longer confined within hadrons [1, 2, 3, 4, 5]. The formation of this strongly interacting plasma is now supported by a broad set of experimental observations. Two of its most compelling signatures are jet quenching [6, 7] and the Number of Constituent Quark (NCQ) scaling of the second-order flow harmonic $v_{2}$ [8, 9]. The observed scaling of elliptic flow of hadrons, collectively follow a universal curve when scaled by their number of constituent quarks, provides direct evidence for the emergence of partonic collectivity prior to hadronization.
Complementary to elliptic flow, the isotropic radial-flow originates from the strong pressure gradients generated at the early stage of the collision, driving the collective transverse expansion of the medium and providing insight into the equation of state and space-time evolution of the QGP [10]. Traditionally, radial flow has been characterized using a transverse momentum ($p_{\rm T}$)-integrated observable, obtained by performing simultaneous blast-wave fits to the $p_{\rm T}$-spectra of identified hadrons. Recently, a new observable, $v_{0}(p_{\mathrm{T}})$, has been introduced to quantify radial flow through the correlation between the particle yield in a given $p_{\mathrm{T}}$-bin and the eventwise mean transverse momentum fluctuation [11, 12]. This differential measure provides a more direct handle on collective expansion and exhibits improved sensitivity compared to traditional blast-wave-based extractions that rely on model-dependent assumptions [13, 14]. Recent measurements by ALICE collaboration [10] on $v_{0}(p_{\mathrm{T}})$ exhibit a clear mass ordering at low $p_{\mathrm{T}}$ and a meson-baryon splitting at intermediate $p_{\mathrm{T}}$, consistent with the qualitative behavior previously observed in $v_{2}$ analyses [15, 16]. The observable is also sensitive to the transport properties of the system, with a pronounced dependence on the bulk viscosity [10, 17], in contrast to elliptic flow, whose magnitude is influenced by both bulk and shear viscous contributions [18, 19]. Inclusive charged-particle measurements demonstrate that the radial-flow observable $v_{0}(p_{\mathrm{T}})$ exhibits genuine collective behavior, manifested through long-range pseudorapidity correlations, factorization of two-particle correlations into single-particle $v_{0}(p_{\mathrm{T}})$, and a nearly centrality-independent normalized shape $v_{0}(p_{\mathrm{T}})/v_{0}$ at low $p_{\mathrm{T}}$ [20, 21]. Further, as reported in Refs. [22, 14], the hydrodynamic-inspired blast-wave model with fluctuating kinetic freeze-out temperature and flow velocity can explain the mass-ordering of $v_{0}(p_{\rm T})$ for identified charged hadrons at low-$p_{\rm T}$. Moreover, as an extension, an AMPT-based study [22] with string melting configuration demonstrated that the meson-baryon crossing at intermediate-$p_{\rm T}$ originates primarily from the quark coalescence mechanism.
A recent investigation of elliptic flow at lower collision energies has established that signatures such as mass ordering can arise even in systems that do not reach full local thermalization. In particular, the study at $\sqrt{s_{NN}}=4.5$ GeV reported in Ref. [9] demonstrated that mass ordering may originate, at least in part, from hadronic interactions during the late stages of the evolution. Such findings introduce an important ambiguity regarding the interpretation of the LHC measurements: the observed collectivity in $v_{0}(p_{\mathrm{T}})$ may receive contributions from both the partonic and hadronic phases.
This motivates a systematic investigation of the dynamical origin of the $v_{0}(p_{\mathrm{T}})$ signal. In the present work, we examine its scaling properties with the number of constituent quarks for the first time within the AMPT String Melting (SM) framework, which provides controlled access to partonic and hadronic transport effects. The results are compared with existing experimental measurements to assess whether the observed behavior can be attributed predominantly to partonic collectivity or can also emerge from hadronic dynamics.
2 Model Description
2.1 The AMPT model
The multi-phase transport (AMPT) model provides a comprehensive microscopic description of heavy-ion collisions and consists of four distinct stages [23]. In the initial stage, HIJING [24, 25] generates the fluctuating nucleon-nucleon interactions, including hard and semi-hard scatterings, jet production, nuclear shadowing, and the associated excited strings. In the default version, these strings hadronize via Lund string fragmentation, whereas in the string-melting (SM) version, they are first converted to their constituent quarks and antiquarks according to the endpoint flavors and momenta, providing a more suitable setup for studying partonic collectivity. The subsequent partonic evolution (present only in the string-melting scenario) is modeled with Zhang’s Parton Cascade (ZPC[26], which includes elastic two-body scatterings among quarks and gluons with parameterized cross-sections, followed by hadronization through a quark coalescence prescription in which nearby partons recombine into mesons and baryons. Finally, the hadronic phase is described by the ART (A Relativistic Transport) model [27], which propagates the produced hadrons through elastic and inelastic meson–meson, meson–baryon, and baryon-baryon interactions, including resonance formation and decay channels. For the present study, we employ the AMPT model version 2.26t9b.
2.2 The PYTHIA8/Angantyr model
The Angantyr model [28] is an extension of the PYTHIA8 framework [29] specifically designed to simulate ppA and AA collisions by stacking multiple fluctuating nucleon-nucleon sub-collisions on top of the standard $pp$ framework, without introducing an explicit hydrodynamic medium or QGP. In Angantyr, the nuclear geometry is determined through a Glauber model calculation that incorporates event-by-event fluctuations of the nucleon–nucleon cross-section. Each binary encounter is subsequently classified as non-diffractive, single-diffractive, or double-diffractive excitation, following a wounded-nucleon-based approach inspired by the Fritiof model [30]. Multi-parton interactions and color reconnection are handled by the underlying PYTHIA8 machinery, which allows Angantyr to generate fully exclusive hadronic final states and to describe global observables such as charged-particle multiplicity and transverse-momentum spectra in heavy-ion collisions [31, 32, 33]. The Angantyr model serves as a non-collective baseline for studies such as radial flow, by quantifying the contributions from initial-state fluctuations and final-state partonic interactions that are not associated with a thermalized medium. The Pb+Pb collision events at $\sqrt{s_{\rm NN}}$ = 5.02 TeV for this analysis are generated using this model.
3 Methodology
3.1 Event Selection and Centrality Determination
In the present analysis, heavy-ion collisions have been simulated for two collision systems. For Au+Au collisions at RHIC energy ($\sqrt{s_{\rm NN}}$ = 200 GeV), a total of 19.9 million minimum-bias events were produced using the AMPT model in its String Melting configuration. For Pb+Pb collisions at LHC energy ($\sqrt{s_{\rm NN}}$ = 5.02 TeV), 3.4 million minimum-bias events were generated using the PYTHIA8/Angantyr model. Centrality, which serves as a proxy for the collision geometry and overlap region, is determined from final-state charged-particle multiplicity. The method is chosen to align with established experimental procedures for each energy regime. In the Au+Au system, centrality for the AMPT-SM events is determined by the charged-particle multiplicity at mid-rapidity ($|\eta|\leq 0.5$), consistent with the standard procedure employed in analyses of RHIC data. For the Pb+Pb system, centrality in the PYTHIA8/Angantyr model is classified using the total charged-particle multiplicity within the pseudorapidity acceptance of the ALICE V0 detector ($2.8<\eta<5.1$ and $3.7<\eta<-1.7$), a method denoted as V0M. This directly mirrors the experimental centrality determination used by ALICE in the measurement of $v_{0}(p_{\rm T})$ [10]. From these samples, specific centrality percentiles are selected for detailed study: 0–20% (most central) and 20–40% (semi-central) for Au+Au collisions, and 10–20%, 30–40%, and 60–70% for Pb+Pb collisions.
@@ 3.1 Event Selection and Centrality Determination
Figure 1: (Color online) Transverse momentum ($p_{\rm T}$) dependence of the radial flow observable $v_{0}(p_{\rm T})$ for identified hadrons in Au+Au collisions generated with the String Melting version of AMPT model, shown for 0-20% (most central) and 20-40% (semi-central) centrality selections. Shaded bands indicate bootstrap-estimated statistical uncertainties.
3.2 Observable
@@
For the present investigation, radial flow is estimated using the novel observable, $v_{0}(p_{\rm T})$, which provides a unique probe of long-range collective dynamics by quantifying event-by-event correlations between fluctuations in charged-particle multiplicity and fluctuations in the average transverse momentum [11]. The observable $v_{0}(p_{\rm T})$ is defined as,
@@
where $N_{\rm ch}^{\rm A}(p_{\rm T})$ is the number of particles in the $p_{\rm T}$ bin and $N_{\rm ch}^{\rm A}$ is the total multiplicity in window A.The second, correlating quantity is the event-averaged mean transverse momentum, measured in a separate, longitudinally-displaced window B (illustrated in Fig. 2), with a gap, $\eta_{\rm gap}>0$ to suppress short-range correlations, is given by,

### v2

# Observation of partonic collectivity via $p_{\rm T}$-differential radial flow fluctuations in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV

@@ Abstract
We report the observation of partonic radial collectivity in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV via the $p_{\rm T}$-differential flow observable $v_{0}(p_{\rm T})$ using the AMPT String Melting model. For inclusive charged hadrons, we establish three signatures of collectivity: long-range pseudorapidity correlations, the factorization of two-particle correlations, and a centrality-independent scaling of $v_{0}(p_{\rm T})$ normalized by its $p_{\rm T}$-integrated value $v_{0}$, analogous to anisotropic flow. For identified particles ($\pi^{\pm},K^{\pm},p+\overline{p}$), the $v_{0}(p_{\rm T})$ spectra show mass ordering at low-$p_{\rm T}$ and meson-baryon separation at intermediate-$p_{\rm T}$. In central collisions, $v_{0}(p_{\rm T})/n_{q}$ exhibits robust Number of Constituent Quark (NCQ) scaling with $(m_{\rm T}-m_{0})/n_{q}$, a scaling that breaks down in peripheral collisions and is more precise at RHIC than at LHC energies, consistent with earlier $v_{2}$ studies. These findings provide strong evidence that radial collectivity originates predominantly at the partonic stage, extending the paradigm of quark-level dynamics from anisotropic to isotropic flow.
I Introduction
Ultra-relativistic nucleus-nucleus collisions provide access to hot and dense QCD matter, commonly known as the quark-gluon plasma (QGP), where quarks are no longer confined within hadrons [30, 24, 11, 6, 9]. The formation of this strongly interacting plasma is now supported by a broad set of experimental observations. Two of its most compelling signatures are jet quenching [19, 27] and the Number of Constituent Quark (NCQ) scaling of the second-order flow harmonic $v_{2}$ [8, 38]. The observed scaling of the elliptic flow of hadrons collectively follows a universal curve when scaled by their number of constituent quarks, providing direct evidence for the emergence of partonic collectivity prior to hadronization.
Complementary to elliptic flow, the isotropic radial flow originates from the strong pressure gradients generated at the early stage of the collision, driving the collective transverse expansion of the medium and providing insight into the equation of state and space-time evolution of the QGP [3]. Traditionally, radial flow has been characterized using a transverse momentum $p_{\rm T}$-integrated observable, obtained by performing simultaneous blast-wave fits to the $p_{\rm T}$-spectra of identified hadrons. Recently, a new observable, $v_{0}(p_{\mathrm{T}})$, has been introduced to quantify radial flow through the correlation between the particle yield in a given $p_{\mathrm{T}}$-bin and the eventwise mean transverse momentum [29, 26]. This differential measure provides a more direct handle on collective expansion and exhibits improved sensitivity compared to traditional blast-wave-based extractions that rely on model-dependent assumptions [33, 28]. Recent measurements of $v_{0}(p_{\mathrm{T}})$ by the ALICE Collaboration [3] show a clear mass-ordering at low-$p_{\mathrm{T}}$ and a meson-baryon splitting at intermediate-$p_{\mathrm{T}}$, along with an approximate NCQ scaling behavior. These features are qualitatively consistent with the systematics previously observed in $v_{2}$ measurements [22, 4, 7, 8, 31], where similar mass-ordering and scaling patterns have been interpreted as signatures of collective expansion and partonic dynamics. Further, the observable is also sensitive to the transport properties of the system, with a pronounced dependence on the bulk viscosity [3, 18], in contrast to elliptic flow, whose magnitude is influenced by both bulk and shear viscous contributions [32, 17]. Inclusive charged-particle measurements by ATLAS collaboration [2] in Pb+Pb collisions demonstrate that the radial-flow observable $v_{0}(p_{\mathrm{T}})$ exhibits genuine collective behavior, manifested through long-range pseudorapidity correlations, factorization of two-particle correlations into single-particle $v_{0}(p_{\mathrm{T}})$, and a nearly centrality-independent normalized shape $v_{0}(p_{\mathrm{T}})/v_{0}$ at low-$p_{\mathrm{T}}$ [2, 12]. Further, as reported in Refs. [35, 28], the hydrodynamic-inspired blast-wave model with fluctuating kinetic freeze-out temperature and flow velocity can explain the mass-ordering of $v_{0}(p_{\rm T})$ for identified charged hadrons at low-$p_{\rm T}$. Moreover, as an extension, an AMPT-based study [35] with string melting configuration demonstrated that the meson-baryon crossing at intermediate-$p_{\rm T}$ originates primarily from the quark coalescence mechanism.
Crucially, the three hallmarks of collectivity and NCQ scaling of $v_{0}(p_{\rm T})$ have thus far been established at LHC energy (Pb+Pb, 5.02 TeV). Whether these signatures persist at RHIC energies remains an open question. In this work, we examine the corresponding factorization properties and scaling behaviors, including the NCQ scaling of $v_{0}(p_{\rm T})$, in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV within the AMPT string-melting framework. Extending these tests to lower collision energies enables an assessment of the energy dependence of the observed collective patterns and provides insight into the robustness of the scaling features reported at the LHC.
The paper is organized as follows. Section II outlines the AMPT and PYTHIA8/Angantyr event generators and the specific configurations used in this study. Section III describes the event selection criteria and the methodology for calculating $v_{0}(p_{\mathrm{T}})$ and its integrated form $v_{0}$. Section IV presents the results, including $v_{0}(p_{\mathrm{T}})$ for inclusive charged hadrons and identified particles. Finally, Section V summarizes the main findings.
II Model Description
II.1 The AMPT model
The multi-phase transport (AMPT) model offers a comprehensive microscopic description of heavy-ion collisions, comprising four distinct stages [23]. In the initial stage, HIJING [36, 20] generates the fluctuating nucleon-nucleon interactions, including hard and semi-hard scatterings, jet production, nuclear shadowing, and the associated excited strings. In the default version, these strings hadronize via Lund string fragmentation, whereas in the string-melting (SM) version, they are first converted to their constituent quarks and antiquarks according to the endpoint flavors and momenta, providing a more suitable setup for studying partonic collectivity. The subsequent partonic evolution (present only in the SM scenario) is modeled with Zhang’s Parton Cascade (ZPC[37], which includes elastic two-body scatterings among quarks and gluons with parameterized cross-sections, followed by hadronization through a quark coalescence prescription in which nearby partons recombine into mesons and baryons. Finally, the hadronic phase is described by the ART (A Relativistic Transport) model [21], which propagates the produced hadrons through elastic and inelastic meson–meson, meson–baryon, and baryon-baryon interactions, including resonance formation and decay channels. For the present study, we employ the AMPT model version 2.26t9b.
II.2 The PYTHIA8/Angantyr model
The Angantyr model [14] extends the PYTHIA8 framework [15] to simulate ppA and AA collisions by superimposing multiple fluctuating nucleon–nucleon sub-collisions within the standard $pp$ formalism, without introducing an explicit hydrodynamic medium or QGP phase. The nuclear geometry is determined via a Glauber-model calculation that includes event-by-event fluctuations of the nucleon–nucleon cross section. Each binary encounter is classified as non-diffractive, single-diffractive, or double-diffractive excitation following a wounded-nucleon-inspired implementation of the Fritiof model [10]. Multi-parton interactions and color reconnection are treated within the underlying PYTHIA8 framework, enabling the generation of fully exclusive hadronic final states and a realistic description of global observables such as charged-particle multiplicity and transverse-momentum spectra in heavy-ion collisions [34, 13, 25]. In this work, PYTHIA8/Angantyr is employed as a non-collective baseline, where particle production arises from superimposed nucleon–nucleon interactions without medium-induced collective expansion. Pb+Pb collision events at $\sqrt{s_{\rm NN}}=5.02$ TeV are generated using version 8.316 of the model.
Figure 1: (Color online) Multiplicity distribution in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV (top); and Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV (bottom), simulated using AMPT String Melting (SM) and PYTHIA8/Angantyr models respectively, with multiple centrality classes from most central to peripheral. The black curve represents the total number of events analyzed.
III Methodology
III.1 Event Selection and Centrality Determination
We analyze Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV and Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. For Au+Au, 36.31 million minimum-bias events are generated using the AMPT model in the string-melting configuration. For Pb+Pb, 5 million minimum-bias events are simulated with PYTHIA8/Angantyr to provide a non-collective baseline. In addition, published ALICE results for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [3] are used for comparison. For Au+Au collisions, centrality is defined via the charged-particle multiplicity at midrapidity ($|\eta|\leq 0.5$). For Pb+Pb collisions, centrality is determined using charged-particle multiplicity in the V0 detector acceptance ($2.8<\eta<5.1$ and $3.7<\eta<-1.7$), corresponding to the V0M estimator employed by ALICE [3]. The multiplicity distributions are shown in Fig. 1, and the corresponding multiplicity intervals for each centrality class are summarized in Tab. 1.
Table 1: Multiplicity ranges ($dN_{\rm ch}/d\eta$) for different centrality intervals in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV (mid-rapidity centrality, AMPT-SM) and Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV (V0M centrality, PYTHIA8/Angantyr).
Au+Au, $\sqrt{s_{\rm NN}}=200$ GeV (AMPT-SM)
Centrality (%) $dN_{\rm ch}/d\eta$
055 527527793793
551010 451451527527
10102020 330330451451
20203030 240240330330
30304040 170170240240
40405050 116116170170
50506060 7676116116
60607070 47477676
70708080 27274747
Pb+Pb, $\sqrt{s_{\rm NN}}=5.02$ TeV (Angantyr)
Centrality (%) $dN_{\rm ch}/d\eta$
10102020 3992399255915591
30304040 1862186227882788
60607070 317317653653
@@ III.1 Event Selection and Centrality Determination
Figure 2: Cartoon illustrating the $\eta$ regions used in the analysis. Regions A (red) and B (green) are separated by a gap $\eta_{\rm gap}$. The observable $F_{\rm A}^{p_{\rm T}}$ is correlated with the transverse momentum $p_{\rm T}]_{\rm B}$ estimated in the opposite region.
@@
Figure 3: (Color online) Transverse momentum dependence of the radial flow observable $v_{0}(p_{\rm T})$ plotted for three $p_{\rm T}^{\rm ref}$ selections at three centrality intervals: 0–5% (most central), 30–40% (semi-central) and 70–80% (peripheral), for Au+Au collisions simulated using AMPT-SM at $\sqrt{s_{\rm NN}}=200$ GeV.
III.2 Observable
@@
For the present investigation, radial flow is estimated using a novel observable, $v_{0}(p_{\rm T})$, which provides a unique probe of long-range collective dynamics by quantifying event-by-event correlations between fluctuations in charged-particle multiplicity and fluctuations in the average transverse momentum [29, 26]. The observable $v_{0}(p_{\rm T})$ is defined as,
@@
where $N_{\rm ch}^{\rm A}(p_{\rm T})$ is the number of particles in the $p_{\rm T}$ bin and $N_{\rm ch}^{\rm A}$ is the total charged-particle multiplicity in window A. The second term in the numerator of Eq. (1) is the event-averaged mean transverse momentum, estimated in a separate, longitudinally-displaced window B (illustrated in Fig. 2), with a gap, $\eta_{\rm gap}>0$ to suppress short-range correlations, is given by,